# Robust Data Pruning: Uncovering and Overcoming Implicit Bias

### v1

@@ 3 Method & Results
We are ready to propose our “fairness-aware” data pruning method, which consists in random subsampling according to carefully selected target class-wise sizes. To imitate the behavior of the cost-sensitive learning procedures that optimize for worst-class accuracy as discussed in Section 2, we propose to select the pruning fraction of each class based on its validation performance given a preliminary model $\psi$ trained on the whole dataset. This query model is still required by all existing data pruning algorithms to compute scores, so we introduce no additional resource overhead.
@@
Input: Target dataset density $d\in[0,1]$. For each class $k\in[K]$: $N_{k}$ samples in the original dataset and validation recall $r_{k}\in[0,1]$.
@@
Algorithm 1 MetriQ
Consider pruning a $K$-way classification dataset originally with $N$ samples down to density $0\leq d\leq 1$, so the target dataset size is $dN$ (prior literature sometimes refers to $1-d$ as the pruning fraction). Likewise, for each class $k\in[K]$, define $N_{k}$ to be the original number of samples so $N=\sum_{k=1}^{K}N_{k}$, and let $0<d_{k}\leq 1$ be the desired density of that class after pruning. Then, we set $d_{k}\propto 1-r_{k}$ where $r_{k}$ denotes recall (accuracy) of class $k$ computed by $\psi$ on a hold-out validation set. In particular, we define Metric-based Quotas (MetriQ) as $d_{k}=d(1-r_{k})/Z$ where $Z=\sum_{k=1}^{K}(1-r_{k})N_{k}\big{/}N$ is a normalizing factor to ensure that the target density is respected, i.e., $dN=\sum_{k=1}^{K}d_{k}N_{k}$. Alas, not all dataset densities $d\in[0,1]$ can be associated with a valid MetriQ collection; indeed, for large enough $d$, the required class proportions may demand $d_{k}>1$ for some $k\in[K]$. In such a case, we do not prune such classes and redistribute the excess density across unsaturated ($d_{k}<1$) classes according to their MetriQ proportions. The full procedure is described in Algorithm 1. This algorithm is guaranteed to terminate as the excess decreases in each iteration of the outer loop.
(a) VGG-16 on CIFAR-10
(b) VGG-19 on CIFAR-100
(c) ResNet-18 on TinyImageNet
Figure 4: The average test performance of various data pruning protocols against dataset density and measures of class-wise fairness. All results averaged over $3$ random seeds. Error bands represent min/max.
Figure 5: Class-wise densities (left) and full-dataset accuracies (right), cf. Figures 2 and 3. Top: VGG-16 on CIFAR-10, Middle: VGG-19 on CIFAR-100, Bottom: ResNet-18 on TinyImageNet.
@@
To validate the effectiveness of random pruning with MetriQ (Random+MetriQ) in reducing classification bias, we compare it to various baselines derived from two of the strongest pruning algorithms: GraNd (Paul et al., 2021), and Forgetting (Toneva et al., 2019). In addition to plain random pruning (Random)—removing a random subset of all training samples—for each of these two strategies, we consider (1) random pruning that respects class-wise ratios automatically determined by the strategy (Random+StrategyQ), (2) applying the strategy for pruning within classes but distributing sample quotas across classes according to MetriQ (Strategy+MetriQ), and (3) the strategy itself. The implementation details and hyperparameter choices are documented in Appendix A.
@@ Results.
As demonstrated in Figure 5, MetriQ produces exceptionally imbalanced datasets unless the density $d$ is too low by heavily pruning easy classes while leaving the more difficult ones intact. As expected from its design, the negative correlation between the full-dataset accuracy and density of each class is much more pronounced for MetriQ compared to existing pruning methods (cf. Figure 3). Based on our examination of these methods in Section 2, we conjectured that these two properties are associated with smaller classification bias, which is well-supported by MetriQ. Not only does it achieve unparalleled performance with random pruning, but it also enhances fairness of GraNd and Forgetting: Strategy+MetriQ curves trace a much better trade-off between the average and worst-class accuracies than their original counterparts (Strategy). At the same time, Random+StrategyQ fares similarly well, surpassing the vanilla algorithms, too. This indicates that robustness is achieved not only from the appropriate class ratios but also from pruning randomly as opposed to cherry-picking difficult samples. We develop more analytical intuition about this phenomenon in Section 4.
Figure 6: The average test performance of various data pruning protocols against dataset density and measures of class-wise fairness. VGG-19 on CIFAR-100. All results averaged over 3 random seeds. Error bands represent min/max.
A fairness baseline.
@@
The results in Figure 4 provide solid evidence that Random+MetriQ is by far the state-of-the-art data pruning algorithm in the fairness framework. Not only is it superior to prior pruning baselines, it in fact produces significantly more robust models compared to the full dataset, too. Thus, we go further and test our method against one cost-sensitive learning technique, Class-wise Difficulty Based Weighted loss (CDB-W) (Sinha et al., 2022). In its simplest form adopted in this study, CDB-W dynamically updates class-specific weights $w_{k,t}=1-r_{k,t}$ at every epoch $t$ by computing recalls $r_{k,t}$ on a holdout validation set, which is precisely the proportions used by MetriQ. Throughout training, CDB-W uses these weights to upweigh the per-sample losses based on the corresponding class labels. As part of the evaluation, we test if the fairness-driven optimization of CDB-W can help data pruning (for these experiments, we use EL2N). To this end, we consider two scenarios: training the final model with CDB-W on the EL2N-pruned dataset (EL2N+CDB-W), and using a fair query model trained by CDB-W to generate EL2N scores (CDB-W+EL2N).
Having access to the full dataset, CDB-W improves the worst-class accuracy of VGG-19 on CIFAR-100 by $5.7\%$ compared to standard optimization, which is almost twice as short of the increase obtained by removing $50\%$ of the dataset with Random+MetriQ (Figure 4). When applied to EL2N-pruned datasets, CDB-W retains that original fairness level across sparsities, which is clearly inferior not only to Random+MetriQ but also to EL2N+MetriQ. Perhaps surprisingly, EL2N with scores computed by a query model trained with CDB-W fails spectacularly, inducing one of the worst bias observed in this study. Thus, MetriQ can compete with other existing methods that directly optimize for worst-class accuracy.
Imbalanced data.
In the literature, fairness issues almost exclusively arise in the context of long-tailed distributions where certain classes or groups appear far less often than others; for example, CDB-W was evaluated in this setting. While dataset imbalance may indeed exacerbate implicit bias of the trained models towards more prevalent classes, our study demonstrates that the key to robustness lies in the appropriate, difficulty-based class proportions rather than class balance per se. Even though the overall dataset size decreases, pruning with MetriQ can produce far more robust models compared to full but balanced datasets (Figure 5). Still, to promote consistency in evaluation and to further validate our algorithm, we consider long-tailed classification scenarios. We follow the approach used by Cui et al. (2019) to inject imbalance into the originally balanced TinyImageNet. In particular, we subsample each class $k\in[K]$ and retain $\mu^{k-1}$ of its original size; for an initially balanced dataset, the size ratio between the largest ($k=1$) and the smallest ($k=K$) classes becomes $\mu^{K-1}$, which is called the imbalance factor denoted by $I$. Figure 7 reveals that Random+MetriQ consistently beats EL2N in terms of both average and robust performance across a range of imbalance factors ($I=2,5,20$). Likewise, it almost always reduces bias of the unpruned imbalanced TinyImageNet even when training with a fairness-aware CDB-W procedure.
@@
Figure 7: The average test performance of Random+MetriQ (red-toned curves) and EL2N (blue-toned curves) against dataset density and measures of class-wise fairness across dataset imbalance factors $I=2,5,20$. ResNet-18 on imbalanced TinyImageNet. All results averaged over 3 random seeds. Error bands represent min/max.
4 Theoretical Analysis
In this section, we derive analytical results for data pruning in a toy model of binary classification for a mixture of two univariate Gaussians with linear classifiers [^1] Our results generalize to isotropic multivariate Gaussians as well, see Appendix B.. Perhaps surprisingly, in this framework we can derive optimal fair class-wise densities that motivate our method, as well as demonstrate how prototypical pruning algorithms fail with respect to worst-class accuracy.
$R(t)=\phi_{0}R_{0}(t)+\phi_{1}R_{1}(t)$ (2)
Figure 8: The effect of different pruning procedures on the solution to the binary classification with the mixture of Gaussians $\mathcal{M}$ defined by $\mu_{0}=-1$, $\mu_{1}=1$, $\sigma_{0}=0.5$, $\sigma_{1}=1$, and $\phi_{0}=\phi_{1}$. Pruning to density $75\%$ (top) and $50\%$ (bottom). Left: Random pruning with the optimal class-wise densities that satisfy $d_{1}\phi_{1}\sigma_{0}=d_{0}\phi_{0}\sigma_{1}$. Middle: SSP. Right: Random pruning with respect to class ratios provided by the SSP algorithm. The overall density of the dataset is $75\%$ (top row) and $50\%$ (bottom row). All results averaged across $10$ datasets $\{D_{i}\}_{i=1}^{10}$ each with $400$ points. The average ERM is $\overline{T}=\frac{1}{10}\sum_{i=1}^{10}T(D^{\prime}_{i})$ fitted to pruned datasets $D^{\prime}_{i}$. The class risks of the average and worst-class optimal decisions for this Gaussian mixture are $R_{0}[t^{*}(1)]=4.8\%$, $R_{1}[t^{*}(1)]=12.1\%$, and $R_{0}(\hat{t})=R_{1}(\hat{t})=9.1\%$.
$\hat{t}=(\mu_{0}\sigma_{1}+\mu_{1}\sigma_{0})/(\sigma_{0}+\sigma_{1}).$ (4)
Note that $\hat{t}<t^{*}(1)$. This means that in the balanced case $\hat{t}$ is closer to $\mu_{0}$, the mean of the “easier” class. To understand how we should prune to achieve optimal worst-class accuracy, we compute class priors $\phi_{0}$ and $\phi_{1}$ that guarantee that the average risk minimization (Equation 2) achieves the best worst-class risk. These priors can be seen as the optimal class proportions in a “fairness-aware” dataset. Observe that, from Equation 3, $t^{*}(\sigma_{0}/\sigma_{1})=\hat{t}$ because the logarithm in the discriminant vanishes and we obtain Equation 4. Therefore, the optimal average risk minimizer coincides with the solution to the worst-case error over classes, and is achieved when the class priors $\phi_{0},\phi_{1}$ satisfy
$\phi_{0}/\phi_{1}=\sigma_{0}/\sigma_{1}.$ (5)
Intuitively, sufficiently large datasets sampled from $\mathcal{M}$ with class-conditionally independent samples mixed in proportions $N_{0}$ and $N_{1}$ should have their empirical 0-1 risk minimizers close to $t^{*}(N_{0}/N_{1})$. That is, randomly subsampling each class $k=0,1$ to densities $d_{k}$ that satisfy $d_{0}N_{0}\sigma_{1}=d_{1}N_{1}\sigma_{0}$ (see Equation 5) yields robust solutions with equal statistical risk across classes and, hence, optimal worst-class error. Figure 8 (left) confirms our analysis: the average empirical risk minimizer (ERM) $\overline{T}$ fitted to datasets pruned in this fashion lands near $\hat{t}$.
The class-conditional independence assumption we made above is crucial. While it is respected when subsampling randomly within each class, it clearly does not hold for existing, more sophisticated, data pruning algorithms. Therefore, even though they tend to prune easier classes more aggressively as evident from Figure 3, they rarely enjoy any improvement of the worst-class performance compared to the original dataset. We further illustrate this observation by replicating a supervised variant of the Self-Supervised Pruning (SSP) developed by Sorscher et al. (2022) for ImageNet. In this algorithm, we remove samples globally (i.e., irrespective of their class membership) located within a certain margin $M>0$ of their class means. Intuitively, this algorithm discards the easiest or the most representative samples from the most densely populated regions of the distribution. As illustrated in Figure 8 (middle), this method fortuitously prunes the easier class more aggressively; indeed, the amount of samples removed from a class with mean $\mu$ and variance $\sigma^{2}$ is proportional to the area under the probability density function over the pruning interval $[\mu-M,\mu+M]$, which is larger for smaller values of $\sigma$. Nevertheless, if the original dataset has classes mixed in proportions $N_{0}$ and $N_{1}$, the solution remains close to $t^{*}(N_{0}/N_{1})$ even after pruning, as we show formally in Appendix B.3. On the other hand, random subsampling according to class-wise pruning proportions defined by SSP does improve the worst-class accuracy, as illustrated in the right plots of Figure 8. This corresponds to our observation in Figure 4 that random pruning respecting the class proportions discovered by GraNd and Forgetting often improves fairness compared to these methods themselves.
Figure 9: (a): Class-wise risk ratios of the optimal solution $t^{*}=t^{*}(\phi_{0}/\phi_{1})$ (MetriQ) vs. optimal ratios based on Equation 5 computed for various $\sigma_{0}$<$\sigma_{1}$ drawn uniformly from $[10^{-2},10^{2}]$ and $\phi_{0}\sim U[0,1]$ and $\phi_{1}=1-\phi_{0}$. The results are independent of $\mu_{0},\mu_{1}$. (b): Random pruning with MetriQ. Left: $d=75\%$; Right: $d=50\%$.
A priori, it is unclear how exactly the variance-based worst-case optimal pruning quotas in Equation 5 generalize to deep learning, and in particular how they can motivate error-based MetriQ, as proposed in Section 3. One could connect our theory to the feature distribution in the penultimate layer of neural networks and determine class densities from cluster variances around class means. In fact, SSP (Sorscher et al., 2022) uses such metrics to determine pruning scores for sample, using a pretrained model. However, such variances are noisy, especially for smaller class sizes, and hard to connect to class accuracies. Therefore, instead, we examine MetriQ in our toy framework. As argued above, given a dataset with class priors $N_{0}$ and $N_{1}$, the optimal class densities must satisfy $d_{0}N_{0}\sigma_{1}=d_{1}N_{1}\sigma_{0}$, while MetriQ requires $d_{0}R_{1}[t^{*}(N_{0}/N_{1})]=d_{1}R_{0}[t^{*}(N_{0}/N_{1})]$.
Figure 9a plots the two proportions for different Gaussian mixtures; we find that MetriQ approximates the optimal quotas quite well, especially when $\sigma_{1}/\sigma_{0}$ is small. Figure 9b demonstrates that random pruning according to MetriQ lands the average ERM near the optimal value. Thus, even though MetriQ does not enjoy simple theoretical guarantees, it still operates near-optimally in this toy setup.
@@ 5 Discussion
Data pruning—removal of uninformative samples from the training dataset—offers much needed efficiency in deep learning. However, all existing pruning algorithms are currently evaluated exclusively on their average performance, ignoring the potential impacts on fairness of the model predictions. Through a systematic study of the classification bias, we reveal that current methods often exacerbate the performance disparity across classes, which can deceptively co-occur with high average performance. At the same time, data pruning arguably operates in a similar manner to techniques that directly optimize for worst-class accuracy, suggesting that there is in fact much potential to improve model fairness by removing the right data. By examining elements of these algorithms, we find evidence that appropriately selecting the class-wise ratios but subsampling randomly within classes should reduce the implicit bias and produce more robust models. This leads us to formulate error-based class-wise pruning quotas coined MetriQ. At the same time, we find value in pruning randomly within classes, as opposed to cherry-picking, which is inherent to the existing data pruning techniques. We confirm the effectiveness of our method on a series of standard computer vision benchmarks; our simple pruning protocol traces the best trade-off between average and worst-class performance among all existing data pruning algorithms and related baselines. Additionally, we find theoretical justification for the phenomenal success of this simple strategy in a toy classification model.
@@
In this study, we focused our empirical evaluation exclusively on classification bias, so we only scratched the surface of fairness in deep learning. Further research is warranted to understand the effect of MetriQ and data pruning in general on worst-group accuracy and spurious correlations.
Acknowledgments and Disclosure of Funding
AV and JK were supported by the NSF Award 1922658. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v2

@@ 4 DRoP: Distributionally Robust Pruning
We are ready to propose our “robustness-aware” data pruning method, which consists in random subsampling according to carefully selected target class-wise sizes, incorporating lessons learned from our theoretical analysis. We propose to select the pruning fraction of each class based on its validation performance given a preliminary model $\psi$ trained on the whole dataset. Such a query model is still required by all existing data pruning algorithms to compute scores, so we introduce no additional resource overhead (see Appendix F for more discussion).
@@
Input: Target dataset density $d\in[0,1]$. For each class $k\in[K]$: original size $N_{k}$, validation recall $r_{k}\in[0,1]$.
@@
Algorithm 1 DRoP
Consider pruning a $K$-way classification dataset originally with $N$ samples down to density $0\leq d\leq 1$, so the target dataset size is $dN$ (prior literature sometimes refers to $1-d$ as the pruning fraction). Likewise, for each class $k\in[K]$, define $N_{k}$ to be the original number of samples so $N=\sum_{k=1}^{K}N_{k}$, and let $0<d_{k}\leq 1$ be the desired density of that class after pruning. Then, we set $d_{k}\propto 1-r_{k}$ where $r_{k}$ denotes recall (accuracy) of class $k$ computed by $\psi$ on a hold-out validation set. In particular, we define DRoP quotas as $d_{k}=d(1-r_{k})/Z$ where $Z=\sum_{k=1}^{K}(1-r_{k})N_{k}\big{/}N$ is a normalizing factor to ensure that the target density is respected, i.e., $dN=\sum_{k=1}^{K}d_{k}N_{k}$. Alas, not all dataset densities $d\in[0,1]$ can be associated with a valid DRoP collection; indeed, for large enough $d$, the required class proportions may demand $d_{k}>1$ for some $k\in[K]$. In such a case, we do not prune such classes and redistribute the excess density across unsaturated ($d_{k}<1$) classes according to their DRoP proportions. The full procedure is described in Algorithm 1. This algorithm is guaranteed to terminate as the excess decreases in each iteration of the outer loop.
@@
To validate the effectiveness of random pruning with DRoP (Random+DRoP) in reducing classification bias, we compare it to various baselines derived from the strongest pruning algorithms: EL2N (Paul et al., 2021) (for CIFAR-10 and ImageNet), GraNd (Paul et al., 2021) and Forgetting (Toneva et al., 2019) (for CIFAR-100 and TinyImageNet). In addition to plain random pruning (Random)—removing a random subset of all training samples—for each of these two strategies, we consider (1) random pruning that respects class-wise ratios automatically determined by the strategy (Random+StrategyQ), (2) applying the strategy for pruning within classes but distributing sample quotas across classes according to DRoP (Strategy+DRoP), and (3) the strategy itself. The motivation for (1) is the anecdotal observation that existing pruning algorithms automatically balance classes according to their validation errors fairly well (Figure 13), while their overly scrupulous filtering within classes may be suboptimal. In support of this view, Ayed & Hayou (2023) formally show that integrating random sampling into score-based pruning procedures improves their average performance. Next, (2) serves as a reasonable ablation, while (3) is the benchmark we compare to. The implementation details are listed in Appendix A.
Figure 5: The average test performance of various data pruning protocols against dataset density and worst-class accuracy. All results averaged over $3$ random seeds. Error bands represent min/max.
@@ Results.
As demonstrated in Figure 6, DRoP produces exceptionally imbalanced datasets unless the density $d$ is too low by heavily pruning easy classes while leaving the more difficult ones intact. As expected from its design, the negative correlation between the full-dataset accuracy and density of each class is much more pronounced for DRoP compared to existing pruning methods (cf. Figure 13).
Figure 6: DRoP. Left: Sorted class densities at different dataset density levels. We report the minimum number of samples per class (SPC) at $10\%$ dataset density. Right: Full dataset test class-wise accuracy against dataset density. We also report the correlation coefficient between these two quantities across classes, averaged over $5$ dataset densities.
@@
Based on our examination of these methods in Section 2, we conjectured that these two properties are associated with smaller classification bias, which is well-supported by DRoP. Not only does it achieve unparalleled performance with random pruning, but it also enhances robustness of GraNd and Forgetting: Strategy+DRoP curves often trace a much better trade-off between the average and worst-class accuracies than their original counterparts (Strategy). At the same time, Random+StrategyQ fares similarly well, surpassing the vanilla algorithms, too. This indicates that robustness is achieved not only from the appropriate class ratios but also from pruning randomly as opposed to cherry-picking hard samples.
A DRO baseline.
@@
The results in Figure 5 provide solid evidence that Random+DRoP is by far the state-of-the-art data pruning algorithm in the robustness framework. Not only is it superior to prior pruning baselines, it in fact produces significantly more robust models compared to the full dataset, too. Thus, we go further and test our method against one representative cost-sensitive learning method from the DRO literature, Class-wise Difficulty Based Weighted loss (CDB-W) (Sinha et al., 2022). In its simplest form adopted in this study, CDB-W dynamically updates class-specific weights $w_{k,t}=1-r_{k,t}$ at every epoch $t$ by computing recalls $r_{k,t}$ on a holdout validation set, which is precisely the proportions used by DRoP. Throughout training, CDB-W uses these weights to upweigh the per-sample losses based on the corresponding class labels. As part of the evaluation, we test if the robustness-driven optimization of CDB-W can help data pruning (for these experiments, we use EL2N). To this end, we consider two scenarios: training the final model with CDB-W on the EL2N-pruned dataset (EL2N+CDB-W), and using a robust query model trained by CDB-W to generate EL2N scores (CDB-W+EL2N).
@@
Figure 7: The average test performance of data pruning protocols against data density and measures of class robustness (VGG-19 on CIFAR-100). All results averaged over 3 random seeds. Error bands represent min/max.
Having access to the full dataset for training, CDB-W improves the worst-class accuracy of VGG-19 on CIFAR-100 by $5.7\%$ compared to standard optimization, which is almost twice as short of the increase obtained by removing $50\%$ of the dataset with Random+DRoP (Figure 5). When applied to EL2N-pruned datasets, CDB-W maintains that original bias across sparsities, which is clearly inferior not only to Random+DRoP but also to EL2N+DRoP. Perhaps surprisingly, EL2N with scores computed by a query model trained with CDB-W fails spectacularly, inducing one of the worst bias observed in this study. Thus, DRoP can compete with other existing methods that directly optimize for worst-class accuracy.
Imbalanced datasets.
In the literature, robustness almost exclusively arises in the context of long-tailed distributions where certain classes or groups appear far less often than others; for example, CDB-W was evaluated in this setting. While dataset imbalance may indeed exacerbate implicit bias of the trained models towards more prevalent classes, our study demonstrates that the key to robustness lies in the appropriate, difficulty-based class proportions rather than class balance per se. Even though the overall dataset size decreases, pruning with DRoP can produce far more robust models compared to full but balanced datasets (Figure 6). Still, to promote consistency in evaluation and to further validate our algorithm, we consider long-tailed classification scenarios. We follow the approach used by Cui et al. (2019) to inject imbalance into the originally balanced TinyImageNet. In particular, we subsample each class $k\in[K]$ and retain $\mu^{k-1}$ of its original size for some $\mu\in(0,1)$. For an initially balanced dataset, the size ratio between the largest ($k=1$) and the smallest ($k=K$) classes then becomes $\mu^{1-K}$, which is called the imbalance factor denoted by $I$. Figure 8 reveals that Random+DRoP consistently beats EL2N in terms of both average and robust performance across a range of imbalance factors ($I=2,5,20$). Likewise, it almost always reduces bias of the unpruned imbalanced TinyImageNet even when training with a robustness-aware CDB-W procedure.
Figure 8: The average test performance of Random+DRoP (red-toned curves) and EL2N (blue-toned curves) against dataset density and measures of class robustness across dataset imbalance factors $I=2,5,20$. ResNet-18 on imbalanced TinyImageNet. Results averaged over 3 random seeds. Error bands represent min/max.
Group Distributional Robustness: Waterbirds.
While we scoped our study around robustness to classification bias, a vast majority of research is concerned with strong model performance across sensitive or minority groups within the distribution. One common benchmark dataset in this area is Waterbirds, which is a binary image classification of bird types where ground truth is spuriously correlated with image background (forest/water) for the majority of samples Sagawa* et al. (2020). Thus, standard optimization techniques produce models that often make predictions based on the background type rather than bird features, yielding poor accuracy for images from waterbird+forest or landbird+water groups. We use the original learning setup introduced by Sagawa* et al. (2020) to evaluate DRoP against other pruning baselines as well as CDB-W. Figure 9 shows that Random+DRoP and EL2N+DRoP achieve a significant improvement of the worst-group accuracy and other metrics. Therefore, as expected, our algorithm applies not only to reduce classification bias but also to improve group-wise robustness.
Figure 9: The average test performance of data pruning protocols against data density and measures of group-wise robustness (ResNet-50 on Waterbirds). Results averaged over 3 random seeds. Error bands represent min/max.
@@ 5 Discussion
Data pruning—removal of uninformative samples from the training dataset—offers much needed efficiency in deep learning. However, all existing pruning algorithms are currently evaluated exclusively on their average performance, ignoring their potentially disparate impact on model predictions across data distribution. Through a systematic study of the classification bias, we reveal that current methods often exacerbate the performance disparity across classes, which can deceptively co-occur with high average performance. This leads us to formulate error-based class-wise pruning quotas coined DRoP. At the same time, we find value in pruning randomly within classes, as opposed to cherry-picking individual samples, which is inherent to the existing data pruning techniques. We confirm the effectiveness of our method on a series of standard computer vision benchmarks; our simple pruning protocol traces the best trade-off between average and worst-class performance among all existing data pruning algorithms and related baselines. Additionally, we find theoretical justification for the phenomenal success of this simple strategy in a toy classification model.
@@
In this study, we focused our empirical evaluation primarily on classification bias. Thus, we only scratched the surface of robustness in deep learning, which is often concerned with group-wise model performance. Further research may attempt to understand the effect of DRoP and data pruning at large on worst-group accuracy and spurious correlations more deeply. Finally, we attribute our contributions mostly to research on data pruning and, therefore, present limited cross-evaluation with a broad spectrum of distributionally robust optimization methods. Thus, future research is warranted to perform a more detailed comparative study between these approaches.